# Predicting Drug-Gene Relations via Analogy Tasks with Word Embeddings

### v2

@@ Introduction
While word embeddings have shown promise in various biomedical applications, their potential for predicting drug-gene relations remains relatively unexplored. To the best of our knowledge, there is limited research specifically focused on utilizing analogy computation of word embeddings to predict drug-gene relations in the biological domain. Our study aims to address this gap by investigating the application of analogy computation of word embeddings to drug-gene relationship prediction.
@@ Methods
Figure 2: 2-d visualization of embeddings using PCA. Drugs and genes are shown in blue and orange, respectively. Solid lines represent the relations. The symbols $\bigstar$ and $\blacksquare$ represent the mean embeddings of the drugs and genes, respectively. The direction of the dashed line connecting these two symbols can be considered as the drug-gene relation vector. (a) Randomly sampled $200$ drug-gene pairs from $\mathcal{R}$. The mean embeddings are computed by Eq. (9). (b) Drugs $d\in\mathcal{D}_{p}$ and genes $g\in\mathcal{G}_{p}$ for $p=$ErbB signaling pathway. Those that have drug-gene relations are shown in boxes. The mean embeddings are computed by Eq. (16).
@@
$\mathrm{E}_{\mathcal{R}}\{\mathbf{u}_{g}\}=\frac{1}{|\mathcal{R}|}\sum_{(d,g)\in\mathcal{R}}\mathbf{u}_{g},\quad\mathrm{E}_{\mathcal{\mathcal{R}}}\{\mathbf{u}_{d}\}=\frac{1}{|\mathcal{R}|}\sum_{(d,g)\in\mathcal{R}}\mathbf{u}_{d}.$ (9)
@@
where $\mathrm{E}_{\mathcal{R}_{p}}\{\cdot\}$ is the sample mean over the set of drug-gene pairs $\mathcal{R}_{p}$. For easy comparison with Eq. (5), Eq. (14) is rewritten as the difference of mean vectors:
$\hat{\mathbf{v}}_{p}=\mathrm{E}_{\mathcal{R}_{p}}\{\mathbf{u}_{g}\}-\mathrm{E}_{\mathcal{R}_{p}}\{\mathbf{u}_{d}\},$ (15)
$\mathrm{E}_{\mathcal{R}_{p}}\{\mathbf{u}_{g}\}=\frac{1}{|\mathcal{R}_{p}|}\sum_{(d,g)\in\mathcal{R}_{p}}\mathbf{u}_{g},\quad\mathrm{E}_{\mathcal{R}_{p}}\{\mathbf{u}_{d}\}=\frac{1}{|\mathcal{R}_{p}|}\sum_{(d,g)\in\mathcal{R}_{p}}\mathbf{u}_{d}.$ (16)
@@
Equation (19) corresponds to Eq. (4). Therefore, similar to the estimator $\hat{\mathbf{v}}$ in Eq. (7), we define an estimator $\hat{\mathbf{v}}^{y}$ for the relation vector $\mathbf{v}^{y}$ as the mean of the vector differences $\mathbf{u}_{g}-\mathbf{u}_{d}$ for $(d,g)\in\mathcal{R}^{y}$:
$\hat{\mathbf{v}}^{y}:=\mathrm{E}_{\mathcal{R}^{y}}\{\mathbf{u}_{g}-\mathbf{u}_{d}\}=\frac{1}{\left|\mathcal{R}^{y}\right|}\sum_{(d,g)\in\mathcal{R}^{y}}(\mathbf{u}_{g}-\mathbf{u}_{d}),$ (20)
where $\mathrm{E}_{\mathcal{R}^{y}}\{\cdot\}$ is the sample mean over the set of drug-gene pairs $\mathcal{R}^{y}$. For easy comparison with Eq. (5), Eq. (20) is rewritten as the difference of mean vectors:
$\hat{\mathbf{v}}^{y}=\mathrm{E}_{\mathcal{R}^{y}}\{\mathbf{u}_{g}\}-\mathrm{E}_{\mathcal{R}^{y}}\{\mathbf{u}_{d}\},$ (21)
$\mathrm{E}_{\mathcal{R}^{y}}\{\mathbf{u}_{g}\}=\quad\frac{1}{|\mathcal{R}^{y}|}\sum_{(d,g)\in\mathcal{R}^{y}}\mathbf{u}_{g},\quad\mathrm{E}_{\mathcal{R}^{y}}\{\mathbf{u}_{d}\}=\quad\frac{1}{|\mathcal{R}^{y}|}\sum_{(d,g)\in\mathcal{R}^{y}}\mathbf{u}_{d}.$ (22)
Given the above, similar to setting G, we perform the analogy tasks in the following setting.
@@
In analogy tasks, we use “known” $\mathcal{R}^{y\mid L_{y}}$ and then predict the target genes $g$ from a drug $d$ for $(d,g)$ in “unknown” $\mathcal{R}^{y\mid U_{y}}$. Using the relation vector $\mathbf{v}^{y\mid L_{y}}$, which represents the drug-gene relations in $\mathcal{R}^{y\mid L_{y}}$, we predict $\mathbf{u}_{g}$ by adding the relation vector $\mathbf{v}^{y}$ to $\mathbf{u}_{d}$:
@@
Equation (24) corresponds to Eq. (19). We define the estimator $\hat{\mathbf{v}}^{y\mid L_{y}}$ for the relation vector $\mathbf{v}^{y\mid L_{y}}$ by using $\mathcal{R}^{y\mid L_{y}}$ instead of $\mathcal{R}^{y}$ in the estimator $\hat{\mathbf{v}}^{y}$ in Eq. (20). Given the above, we perform the analogy tasks in the following setting.
@@
BioConceptVec [13] provides four pretrained 100-dimensional word embeddings: CBOW [7], skip-gram [7, 8], GloVe [33], and fastText [34]. Since CBOW is a simpler model than skip-gram and skip-gram performs better in analogy tasks compared to GloVe [9], we used BioConceptVec skip-gram embeddings for our experiments. Note that fastText can essentially be considered as a skip-gram using $n$-grams. To complement the pretrained BioConceptVec embeddings, we further trained 300-dimensional skip-gram embeddings on the publicly available PubMed abstracts. To train our embeddings, we used PubTator [14] to normalize six major biological concepts (genes, mutations, diseases, chemicals, cell lines, and species) in PubMed abstracts, and then tokenized them using NLTK [35]. Note that since widely used embeddings such as word2vec and GloVe typically have 300 dimensions, we set the dimensions of our embeddings to 300 instead of the original 100. The hyperparameters used to train our embeddings are shown in Table S4 in Supplementary Information 1.4.
For BioConceptVec and our skip-gram embeddings, Table 2 shows some basic statistics for setting G. Due to the difference in training data size and minimal word occurrence, the sizes of certain sets such as $|\mathcal{D}|$ and $|\mathcal{G}|$ differ significantly between BioConceptVec and our skip-gram embeddings, but the sizes of other sets show somewhat similar trends. Table S1 in Supplementary Information 1.1.1 shows the statistics for settings P1 and P2. The hyperparameters used to train our skip-gram are shown in Table S4 in Supplementary Information 1.4.
@@
To evaluate the performance of predicting target genes by adding the relation vector to a drug embedding, we compare it to baseline methods. As a simple baseline, we randomly select one gene from the search space for each query. We refer to this baseline as the random baseline. In the random baseline, we repeated the experiments 10 times and used the average score as the final result. In addition to this, as strong generative model baselines, we used the Generative Pre-trained Transformer (GPT) [37] series to predict the top 10 target genes from a query drug by zero-shot. We used the GPT-3.5 [38], GPT-4 [39], and GPT-4o (https://platform.openai.com/docs/models/gpt-4o) models with the temperature hyperparameter set to 00. See Supplementary Information 1.6 for more details on the models and prompt template for the predictions.
@@
As evaluation metrics, we use the top-$k$ accuracy explained in each setting, especially the top-1 and top-10 accuracies. We also use Mean Reciprocal Rank (MRR) as another evaluation metric. MRR is a statistical measure to evaluate search performance, expressed as the sum of the reciprocals of the ranks at which the predicted results first appear in the answer set. Note that since MRR cannot be calculated for the GPT models due to the limited number of prediction candidates output by GPT, only top-1 and top-10 accuracies are calculated.
@@ Results
BioConceptVec G Random 0.000 0.000 0.000
$\hat{\mathbf{v}}$ 0.304 0.645 0.419
P1 Random 0.000 0.000 0.000
$\hat{\mathbf{v}}_{p}$ 0.499 0.790 0.602
P2 Random 0.000 0.000 0.000
$\hat{\mathbf{v}}_{p}$ 0.522 0.807 0.624
Our embeddings G Random 0.000 0.000 0.001
$\hat{\mathbf{v}}$ 0.300 0.686 0.426
P1 Random 0.000 0.001 0.001
$\hat{\mathbf{v}}_{p}$ 0.589 0.862 0.685
P2 Random 0.000 0.001 0.001
$\hat{\mathbf{v}}_{p}$ 0.600 0.880 0.700
@@
Table 3: Gene prediction performance in settings G, P1, and P2.
@@
The random baseline in settings G, P1, and P2 showed nearly zero scores for all evaluation metrics, suggesting that random prediction is nearly impossible in these analogy tasks. This also confirms the superior performance of prediction by adding relation vectors. Among the GPT series, GPT-4o performed the best, followed by GPT-4, which outperformed GPT-3.5. In top-1 accuracy, the GPT series performed better than predictions made by adding relation vectors. However, in top-10 accuracy, the difference becomes smaller. Notably, predictions using $\mathbf{v}_{p}$ with our skip-gram embeddings in settings P1 and P2 performed comparably to GPT-4o and outperformed both GPT-3.5 and GPT-4.
@@
In setting Y1, the number of queries increases steadily from older to more recent years, while in setting Y2, the number of queries increases initially and then decreases over time. In addition, the size of the search space is small for older years (e.g., $|\mathcal{G}^{1975}|=725$) and increases for later years (e.g., $|\mathcal{G}^{2020}|=47509$). As a result, in both settings Y1 and Y2, the datasets for older years such as $y=1975$, and $1980$ have fewer queries and smaller search spaces, resulting in unusually high evaluation metric scores. In setting Y2, the datasets for later years have fewer queries and larger search spaces, which tends to produce lower evaluation metric scores. In both settings, the evaluation metric scores for the random baseline are close to zero. For more detailed results for settings Y1 and Y2, see Table S12 in Supplementary Information 2.2.1.
@@
The analogy tasks in settings P1Y1 and P1Y2 use detailed pathway information. Therefore, predicting the relations that appeared after year $y$ is easier in settings P1Y1 and P1Y2 than in settings Y1 and Y2. In addition, the results for the older years show high evaluation metric scores in settings P1Y1 and P1Y2, similar to those in settings Y1 and Y2. The evaluation metric scores of the random baseline are close to zero.
@@ Discussion
Mikolov et al. [54] has shown that skip-gram achieves a top-1 accuracy of 0.56 on the Microsoft Research Syntactic Analogies Dataset [55]. Given this, the performance of BioConceptVec and our skip-gram embeddings in the analogy tasks for drug-gene pairs, as presented in Table 3, is comparable to that of skip-gram in the basic analogy tasks. Also, Tshitoyan et al. [18] reported an overall accuracy of 60.1% when using analogy computation of word embedding to predict 29,046 relationships across various concepts in materials science, including element names, crystal symmetries, and magnetic properties. Our results for drug-gene relationship prediction show a similar level of accuracy, suggesting that the approach of using analogy computation of word embedding can be effective in the biomedical domain as well. This demonstrates the potential of this method across different scientific disciplines. However, it’s important to note that the accuracy varies significantly across different types of relationships in the materials science study. For instance, predictions for chemical element names showed high accuracy (71.4%), while crystal structure names had lower accuracy (18.7%). This variability suggests that the effectiveness of word embedding analogies may depend on the specific type of relationship being predicted.
@@
In Table 3, predictions using $\mathbf{v}_{p}$ with our skip-gram embeddings in settings P1 and P2 performed on par with GPT-4o and outperformed both GPT-3.5 and GPT-4. It is rather surprising that simple vector addition can achieve performance comparable to large language models. Given that the GPT series is trained on vast amounts of text data, using relation vector addition is an efficient method. Furthermore, the results showing that target genes can be predicted with performance comparable to or better than large language models imply that high-quality information about drug-gene relationships is encoded in the embeddings. This also suggests that the computation of analogies may be a fundamental inference principle within large language models.
@@
In addition, we divided the datasets by year and redefined the vectors representing known relations to predict unknown relations. The experimental results presented in Fig. 3 demonstrate that our approach can predict unknown future relations to some extent. The experiments conducted with year-split datasets are a novel attempt not found in existing research, and confirmed the effectiveness of predicting unknown future events.

### v4

@@ Introduction
While word embeddings have shown promise in various biomedical applications, their potential for predicting drug-gene relations remains relatively unexplored. To the best of our knowledge, there is limited research specifically focused on utilizing analogy computation of word embeddings to predict drug-gene relations in the biological domain. Our study aims to address this gap by investigating the application of analogy computation of word embeddings to drug-gene relation prediction.
@@ Methods
Figure 2: 2-d visualization of embeddings using PCA. Drugs and genes are shown in blue and orange, respectively. Solid lines represent the relations. The symbols $\bigstar$ and $\blacksquare$ represent the mean embeddings of the drugs and genes, respectively. The direction of the dashed line connecting these two symbols can be considered as the drug-gene relation vector. (a) Randomly sampled $200$ drug-gene pairs from $\mathcal{R}$. The mean embeddings are computed by Eq. (9). (b) Drugs $d\in\mathcal{D}_{p}$ and genes $g\in\mathcal{G}_{p}$ for $p=$ErbB signaling pathway. Those that have drug-gene relations are shown in boxes. The mean embeddings are computed by Eq. (S2) in Supplementary Information1.1.1.
@@
$\mathrm{E}_{\mathcal{\mathcal{R}}}\{\mathbf{u}_{d}\}=\frac{1}{|\mathcal{R}|}\sum_{(d,g)\in\mathcal{R}}\mathbf{u}_{d},\quad\mathrm{E}_{\mathcal{R}}\{\mathbf{u}_{g}\}=\frac{1}{|\mathcal{R}|}\sum_{(d,g)\in\mathcal{R}}\mathbf{u}_{g}.$ (9)
@@
where $\mathrm{E}_{\mathcal{R}_{p}}\{\cdot\}$ is the sample mean over the set of drug-gene pairs $\mathcal{R}_{p}$.
@@
Equation (17) corresponds to Eq. (4). Therefore, we define the estimator $\hat{\mathbf{v}}^{y}$ for the relation vector $\mathbf{v}^{y}$ by using $\mathcal{R}^{y}$ instead of $\mathcal{R}$ in the estimator $\hat{\mathbf{v}}$ in Eq. (7). Given the above, similar to setting G, we perform the analogy tasks in the following setting.
@@
In analogy tasks, we use “known” $\mathcal{R}^{y\mid L_{y}}$ and then predict the target genes $g$ from a drug $d$ for $(d,g)$ in “unknown” $\mathcal{R}^{y\mid U_{y}}$. Using the relation vector $\mathbf{v}^{y\mid L_{y}}$, which represents the drug-gene relations in $\mathcal{R}^{y\mid L_{y}}$, we predict $\mathbf{u}_{g}$ by adding the relation vector $\mathbf{v}^{y\mid L_{y}}$ to $\mathbf{u}_{d}$:
@@
Equation (19) corresponds to Eq. (17). Therefore, we define the estimator $\hat{\mathbf{v}}^{y\mid L_{y}}$ for the relation vector $\mathbf{v}^{y\mid L_{y}}$ by using $\mathcal{R}^{y\mid L_{y}}$ instead of $\mathcal{R}$ in the estimator $\hat{\mathbf{v}}$ in Eq. (7). Given the above, we perform the analogy tasks in the following setting.
@@
BioConceptVec [13] provides four pretrained 100-dimensional word embeddings: CBOW [7], skip-gram [7, 8], GloVe [33], and fastText [34]. Since CBOW is a simpler model than skip-gram and skip-gram performs better in analogy tasks compared to GloVe [9], we used BioConceptVec skip-gram embeddings for our experiments. Note that fastText can essentially be considered as a skip-gram using $n$-grams. To complement the pretrained BioConceptVec embeddings, we further trained 300-dimensional skip-gram embeddings on the publicly available PubMed abstracts. As with the original BioConceptVec, to train our embeddings, we used PubTator [14] to convert six major biological concepts (genes, mutations, diseases, chemicals, cell lines, and species) in PubMed abstracts into their respective IDs, followed by tokenization using NLTK [35]. Note that since widely used embeddings such as word2vec and GloVe typically have 300 dimensions, we set the dimensions of our embeddings to 300 instead of the original 100. The hyperparameters used to train our embeddings are shown in Table S4 in Supplementary Information 1.4.
For BioConceptVec and our skip-gram embeddings, Table 2 shows some basic statistics for setting G. Due to the difference in training data size and minimal word occurrence, the sizes of certain sets such as $|\mathcal{D}|$ and $|\mathcal{G}|$ differ significantly between BioConceptVec and our skip-gram embeddings, but the sizes of other sets, such as $|\mathcal{R}|$, which represents the size of overall relations, show somewhat similar trends. Figure S2 in Supplementary Information 1.4 shows the distribution of the sizes of the answer sets for each drug $d$ under settings G, P1, and P2 for both BioConceptVec and our skip-gram embeddings. Table S1 in Supplementary Information 1.1.1 shows the statistics for settings P1 and P2. The hyperparameters used to train our skip-gram are shown in Table S4 in Supplementary Information 1.4.
@@
To evaluate the performance of predicting target genes by adding the relation vector to a drug embedding, we compare it to baseline methods.
As a simple yet reasonable baseline, we randomly sampled genes from the set $G$ containing the genes that have drug-gene relations with at least one drug. To increase the probability of sampling genes with more drug-gene relations, a gene $g\in G$ was sampled with a probability proportional to $|[g]|$, the size of the set of drugs that have correct drug-gene relations with $g$. In other words, sampling was performed with probabilities proportional to the number of queries that have correct relations. For simplicity, this sampling method was consistently used across settings G, P1, and P2. We refer to this baseline as the random baseline. In this baseline, we repeated the experiments 10 times and used the average score as the final result. Similar random baselines can also be applied to settings by year and to settings for predicting drugs from genes. See Supplementary Information 1.6 for details.
The analogy task we consider can be regarded as a task of predicting relationships between entities. Therefore, we adopted Knowledge Graph Embedding (KGE) [37] as a general baseline method for predicting drug-gene relations. KGE learns embeddings for head entities, tail entities, and relations from triplets by leveraging the explicit structure of the knowledge graph. Using the learned embeddings, it predicts a specific tail corresponding to a given head and relation. Among various KGE methods, we used TransE [38], one of the most representative approaches, with the embedding dimensions set to the commonly used size of 500. Similar to the analogy computation, TransE predicts the tail embedding by adding the relation vector to the head embedding. However, while the analogy computation relies solely on document data and pathway information, TransE leverages explicit relations in the knowledge graph to directly learn embeddings and relation vectors. Word embeddings such as skip-gram are learned from large corpora (e.g., PubMed abstracts) without being tied to any particular research field or specific target task. Consequently, they capture broad linguistic and semantic information (not restricted to drug-gene relations), enabling them to be applied to various tasks (e.g., analogy tasks [39], sentence similarity tasks [5]). In contrast, KGE is learned based on knowledge and relational information tailored to a specific task, resulting in embeddings specialized for that task. For details on the dataset splits and hyperparameters used for training TransE, see Supplementary Information 1.8.
In addition to this, as strong generative model baselines, we used the Generative Pre-trained Transformer (GPT) [40] series to predict the top 10 target genes from a query drug by zero-shot. We used the GPT-3.5 [41], GPT-4 [42], and GPT-4o (https://platform.openai.com/docs/models/gpt-4o) models with the temperature hyperparameter set to 00. See Supplementary Information 1.7 for more details on the models and prompt template for the predictions.
@@
As evaluation metrics, we use the top-$k$ accuracy explained in each setting, especially the top-1 and top-10 accuracies. We also use Mean Reciprocal Rank (MRR) as another evaluation metric. MRR is a statistical measure to evaluate search performance, expressed as the sum of the reciprocals of the ranks at which the predicted results first appear in the answer set. Note that, in the random baseline, all genes $g\in G$ are ranked without duplication through sampling with probabilities proportional to $|[g]|$. By treating these rankings as prediction results, we can calculate not only top-1 and top-10 accuracies but also MRR. Since MRR cannot be calculated for the GPT models due to the limited number of prediction candidates output by GPT, only top-1 and top-10 accuracies are calculated.
@@ Results
BioConceptVec G Random 0.019 0.137 0.063
TransE 0.576 0.682 0.615
analogy 0.304 0.645 0.419
P1 Random 0.018 0.140 0.063
TransE 0.745 0.871 0.792
analogy 0.499 0.790 0.602
P2 Random 0.022 0.162 0.072
TransE 0.813 0.896 0.842
analogy 0.522 0.807 0.624
Our embeddings G Random 0.020 0.140 0.065
TransE 0.571 0.689 0.614
analogy 0.300 0.686 0.426
P1 Random 0.020 0.142 0.066
TransE 0.752 0.873 0.798
analogy 0.589 0.862 0.685
P2 Random 0.023 0.164 0.074
TransE 0.819 0.907 0.854
analogy 0.600 0.880 0.700
@@
Table 3: Gene prediction performance in settings G, P1, and P2. In analogy tasks, the relation vector for prediction is $\hat{\mathbf{v}}$ in setting G and $\hat{\mathbf{v}}_{p}$ in settings P1 and P2. In the random baseline, a gene $g\in G$ is sampled with probabilities proportional to the size of the set of drugs related to $g$. TransE is evaluated using a different data split, while for the GPT series, only the best scores across settings are shown in the table. Since TransE’s performance depends on the size of the training data, Fig. 4 compares analogy computation with TransE under varying proportions of the training data.
@@
The performance of the random baseline in settings G, P1, and P2 was consistently low across all evaluation metrics. Although random sampling was restricted to genes in the set $G$, these results demonstrate that random prediction is challenging in these analogy tasks. This also confirms the superior performance of prediction by adding relation vectors. TransE, as a KGE method, naturally outperforms our proposed approach across all three evaluation metrics. Interestingly, however, the top-10 accuracy of the prediction by vector addition is comparable to that of TransE. This suggests that our method has the latent potential to achieve performance close to TransE without directly exploiting explicit relational information. We then hypothesize that the strong performance of TransE observed in Table 3 depends on the size of the training data. In the Discussion section, we compare analogy computation and TransE under varying proportions of the training data (Fig. 4). Among the GPT series, GPT-4o performed the best, followed by GPT-4, which outperformed GPT-3.5. In top-1 accuracy, the GPT series performed better than predictions made by adding relation vectors. However, in top-10 accuracy, the difference becomes smaller. Notably, predictions using $\mathbf{v}_{p}$ with our skip-gram embeddings in settings P1 and P2 performed comparably to GPT-4o and outperformed both GPT-3.5 and GPT-4.
@@
In setting Y1, the number of queries increases steadily from older to more recent years, while in setting Y2, the number of queries increases initially and then decreases over time. In addition, the size of the search space is small for older years (e.g., $|\mathcal{G}^{1975}|=725$) and increases for later years (e.g., $|\mathcal{G}^{2020}|=47509$). As a result, in both settings Y1 and Y2, the datasets for older years such as $y=1975$, and $1980$ have fewer queries and smaller search spaces, resulting in unusually high evaluation metric scores. In setting Y2, the datasets for later years have fewer queries and larger search spaces, which tends to produce lower evaluation metric scores. In both settings, the performance of the random baseline is extremely high for older years such as $y=1975$ and $y=1980$, where the number of genes related to drugs is small. However, after $y=1985$, our method consistently outperforms the random baseline. For more detailed results for settings Y1 and Y2, see Table S12 in Supplementary Information 2.2.1.
@@
The analogy tasks in settings P1Y1 and P1Y2 use detailed pathway information. Therefore, predicting the relations that appeared after year $y$ is easier in settings P1Y1 and P1Y2 than in settings Y1 and Y2. In addition, the results for the older years show high evaluation metric scores in settings P1Y1 and P1Y2, similar to those in settings Y1 and Y2. The performance of the random baseline also shows similar trends to those observed in settings Y1 and Y2.
@@ Discussion
Mikolov et al. [57] has shown that skip-gram achieves a top-1 accuracy of 0.56 on the Microsoft Research Syntactic Analogies Dataset [39]. Given this, the performance of BioConceptVec and our skip-gram embeddings in the analogy tasks for drug-gene pairs, as presented in Table 3, is comparable to that of skip-gram in the basic analogy tasks. Also, Tshitoyan et al. [18] reported an overall accuracy of 60.1% when using analogy computation of word embedding to predict 29,046 relationships across various concepts in materials science, including element names, crystal symmetries, and magnetic properties. Our results for drug-gene relation prediction show a similar level of accuracy, suggesting that the approach of using analogy computation of word embedding can be effective in the biomedical domain as well. This demonstrates the potential of this method across different scientific disciplines. However, it’s important to note that the accuracy varies significantly across different types of relationships in the materials science study. For instance, predictions for chemical element names showed high accuracy (71.4%), while crystal structure names had lower accuracy (18.7%). This variability suggests that the effectiveness of word embedding analogies may depend on the specific type of relationship being predicted.
Strengths and insights from prediction by adding the relation vector
Figure 4: Changes in evaluation metrics on the test data for analogy computation and TransE under setting P1 as the proportion of training data varies. Analogy computation maintains consistent performance regardless of the proportion of training data, and outperforms TransE when the proportion of training data is small. See Supplementary Information 2.4 for experimental details.
To test the hypothesis that the superior performance of TransE depends on the size of the training data, we compared the performance of our method with that of TransE in setting P1 by varying the proportion of training data used to compute relation vectors (see Supplementary Information 2.4 for experimental details). As Fig. 4 shows, TransE’s performance degrades as the proportion of training data decreases, whereas our method maintains nearly constant performance. In particular, our method outperforms TransE when the available training data is small. This is because our approach derives relation vectors from already-trained word embeddings, while TransE learns all embeddings from scratch using the training data. These findings suggest that our method can function efficiently even with limited datasets.
@@
In addition, in Table 3, predictions using $\mathbf{v}_{p}$ with our skip-gram embeddings in settings P1 and P2 performed on par with GPT-4o and outperformed both GPT-3.5 and GPT-4. It is rather surprising that simple vector addition can achieve performance comparable to large language models. Given that the GPT series is trained on vast amounts of text data, using relation vector addition is an efficient method. Furthermore, the results showing that target genes can be predicted with performance comparable to or better than large language models imply that high-quality information about drug-gene relations is encoded in the embeddings. This also suggests that the computation of analogies may be a fundamental inference principle within large language models.
@@
Furthermore, we divided the datasets by year and redefined the vectors representing known relations to predict unknown relations. The experimental results presented in Fig. 3 demonstrate that our approach can predict unknown future relations to some extent. The experiments conducted with year-split datasets are a novel attempt not found in existing research, and confirmed the effectiveness of predicting unknown future events.
Biological information intrinsic to the embedding space
@@
Limitations
This study has several limitations.
First, we focus exclusively on drug-gene relations. Consequently, the applicability of this method to other biological relationships remains untested.
Additionally, as shown in Table 3, our approach for predicting genes from drugs by simply adding relation vectors performs worse than KGE models such as TransE in terms of accuracy. While the goal of this study is not to improve performance but to explore the information intrinsic to the embedding space, efforts to improve performance remain an important future direction.
Furthermore, the embeddings used in this study are static embeddings learned with skip-gram, which cannot handle out-of-vocabulary concepts. Assigning a single embedding to a concept that serves different roles depending on the context (e.g., concepts used differently in the past and present) may also be suboptimal.
Future Directions
Building on the limitations of this study, several future directions can be considered.
First, since BioConceptVec includes concepts such as diseases and mutations, it would be useful to investigate whether analogy computations using relation vectors can also be applied to relationships beyond drugs and genes. This could serve as a valuable follow-up study.
Additionally, using dynamic embeddings instead of static embeddings represents another promising direction. Since models like BERT [4] compute embeddings based on context, taking advantage of this feature may allow more accurate modeling of drug-gene relations. For example, this approach could help capture relations that vary with context, as well as drugs whose applications evolve over time.
Furthermore, it would be interesting to investigate how the properties of drugs or genes affect performance in analogy tasks. For example, if the drug-gene relations of a drug have been extensively studied, can related genes for that drug be more easily predicted? To address this question, we conducted a simple experiment to examine the correlation between the size of the answer set for each drug and the rank of the search results predicted by adding relation vectors. Scatter plots of these values for settings G, P1, and P2, using BioConceptVec and our embeddings, are shown in Fig. S5 in Supplementary Information 2.3. Mathematically, a larger answer set is expected to make prediction easier, resulting in search result ranks closer to 1. The results showed a slight negative correlation in setting G, as expected (BioConceptVec: $-0.201$, our embeddings: $-0.196$). However, in settings P1 and P2, little to no correlation was observed (BioConceptVec: $0.054$ and $-0.102$, our embeddings: $0.079$ and $-0.041$). This suggests that in settings P1 and P2, incorporating pathway information into relation vectors helps maintain good search results even when the answer sets are small. By performing such detailed analyses, we can further deepen our understanding of the biological knowledge intrinsic to the embedding space.